OpenStudy (chunkymonkay):

Which answer shows this equation in standard form? 7-3(x-y)=-5x+2 A. -8x+3y=-5 B. -2x-3y=5 C.2x+3y=-5 D.2x-3y=-5

OpenStudy (danjs):

simplify, First distribute the -3 into the parenthesis ....A*(B+C) = A*B+A*C

OpenStudy (chunkymonkay):

is the answer d cuz that's what I think it is

OpenStudy (danjs):

k, show me the work you did to get there , ill check it out

OpenStudy (danjs):

After the 3 is distributed into the parenthesis you get 7 - 3x + 3y = -5x + 2

OpenStudy (chunkymonkay):

im still a little bit confused

OpenStudy (chunkymonkay):

@Godlovesme

OpenStudy (godlovesme):

did u distribute -3 like @DanJS told u?

OpenStudy (chunkymonkay):

no I don't know how :( would it be 7-3=-5x+2?

OpenStudy (godlovesme):

 so far so good?

OpenStudy (chunkymonkay):

oh okay than whats next?

OpenStudy (godlovesme):



OpenStudy (chunkymonkay):

thank you :)

OpenStudy (godlovesme):

welcome:D